# Contrastive Language–Structure Pre-training Driven by Materials Science Literature

### v1

@@ 1 Introduction
AI-driven materials science pursues this ambition through the use of machine learning (ML). One area of research has focused on predicting material properties using graph neural networks [3, 4, 5, 6] and transformers [7, 8], leveraging large-scale crystal structure datasets annotated with properties simulated by first-principles calculations. Although this approach has shown success, these models are specialized for specific simulatable properties, such as bandgap, and are unable to provide a comprehensive view of materials with diverse properties and functionalities.
Other studies have explored developing embedding spaces for crystal structures to capture their similarities based on properties and functionalities [9, 10, 11, 12]. However, these efforts are bottlenecked by the lack of dedicated training datasets with diverse property and functionality annotations. Annotating crystal structures is costly, requiring expert knowledge, physical experimentation, or computationally intensive simulations. Consequently, these methods often produce abstract, unannotated embedding spaces that are not easily navigable for materials discovery.
@@
To overcome this limitation, we introduce Contrastive Language–Structure Pre-training (CLaSP) (Fig. 1). CLaSP leverages a large-scale dataset of published crystal structures and corresponding article information retrieved from the Crystallography Open Database (COD) [22]. Specifically, we utilize publication titles for pre-training and keyword-based captions, generated from pairs of titles and abstracts using a large language model (LLM), for fine-tuning.
We hypothesize that these textual sources provide a comprehensive representation of material characteristics. By leveraging bibliographic data as natural descriptions of structures, we bypass the need for labor-intensive, specialized annotations, enabling large-scale training of crossmodal models. The effectiveness of CLaSP is demonstrated through two key applications: intuitive text-based material retrieval and materials space mapping. Extensive analyses show that CLaSP effectively learns structure embeddings that capture abstract and complex material concepts, such as ‘superconductor’ and ‘metal-organic frameworks’.
@@ 2 Contrastive language–structure pre-training
To learn crystal embeddings $\{\bm{c}_{i}\}$, a CGCNN model [3] was trained from scratch to serve as the crysta encoder. For text embeddings $\{\bm{t}_{i}\}$, a frozen pre-trained SciBERT model [24] was used as the text encoder, followed by a three-layer multilayer perceptron (MLP) fed with the CLS token embedding. We used the embedding dimensionality of 768 for both modalities. We provide the detailed training procedure in Sec. 7.1.
@@ 3.1 Crystal structure modeling
Crystal structure modeling using neural networks serves as a fundamental basis for ML models that process crystal structures. One of its most straightforward applications is crystal property prediction, for which various crystal encoder architectures have been proposed. These architectures typically employ geometric graph neural networks (GNNs) that model interatomic interactions via neural message passing. Early successful examples include CGCNN [3] and SchNet [30], which represent crystal structures as distance-based graphs. Subsequent work has aimed to capture richer interaction representations by incorporating bond and global state attributes [4], three-body interactions [31], attention mechanisms [32, 7, 8], and infinite periodicity [6, 8].
@@ 3.2 Embedding learning for materials and crystal structures
More recent studies have explored extending material representations through multimodal learning with texts [20, 21], aiming to attribute semantic meaning to material embeddings. For example, Ozawa et al. [21] used crystal structures along with rule-based textual descriptions of their geometric features for contrastive learning. However, these methods rely on textual descriptions of structural features rather than material properties, thus limiting their ability to capture high-level information such as material functionalities.
@@ 4 Results
In this section, we demonstrate the effectiveness of the proposed approach both quantitatively and qualitatively through two applications: text-based material retrieval and embedding space visualization. For text-based material retrieval (Sec. 4.1), we evaluate how accurately the model links crystal structures with their textual descriptions, such as ‘superconductor’ and ‘narrow-bandgap material’. Additionally, we verify the design of our loss function through this retrieval application (Sec. 4.2). For embedding space visualization, we construct an intuitive map of the materials space (Sec. 4.3) and compare it with an existing embedding approach (Sec. 4.4).
@@ 4.1 Zero-shot crystal structure screening by texts
To evaluate CLaSP’s ability to link crystal structures with textual property descriptions, we performed crystal structure retrieval using keywords representing material functionalities (e.g., ‘thermoelectric’ and ‘superconductor’). Given the embedding of a query keyword, we retrieved structure embeddings from the test set that showed high cosine similarities with the keyword embedding. We regarded a structure to possess a queried property if the associated paper title contained the keyword or its variations (e.g., given ‘superconductor’ as a query, the terms ‘superconductive’ and ‘superconductivity’ were also considered correct).
We evaluated the trade-off between true positives and false positives using the ROC (receiver operatorating characteristic) and ROC-AUC (area under the ROC curve). Additionally, we assessed Average Precision (AP). Given the extremely low positive-to-negative ratios in our retrieval tasks (i.e., 20 to 165 positives in a total of 40,604 samples for each keyword), we evaluated AP using a balanced random subset of negative samples to make the AP scores more informative.
Figure 3 shows ROC curves for six query keywords before and after fine-tuning, with dashed diagonal lines representing random selection. For a more comprehensive evaluation, Table 1 summarizes the mean and individual ROC-AUC scores for the six keywords, along with the Average Precision (AP) scores. While the zero-shot prediction using the pre-trained model (Fig. 3, left) already demonstrates good performance, with a mean ROC-AUC of 0.7185, fine-tuning (Fig. 3, right) further improved it to 0.7804. A similar trend is observed in the AP evaluation (Table 1). These results highlight the ability of the CLaSP model to capture complex structure–property relationships across diverse material functionalities by analyzing crystal structures alone. Section 5 will discuss these findings in more detail.
@@
Figure 3: ROC curves of keyword-based crystal structure retrieval. The zero-shot results with only pre-training (left) show good performance, and fine-tuning leads to further improvements (right). The test set consists of 40,604 materials, and the dashed diagonal lines represent the ROC curves for random selection.
@@
ROC-AUC AP
Query # Positives PT FT PT FT
ferromagnetic 165 0.4281 0.5674 0.5018 0.5716
ferroelectric 95 0.5814 0.7097 0.5357 0.6380
semiconductor 90 0.7382 0.8290 0.7705 0.8254
electroluminescence 24 0.8106 0.8090 0.9050 0.8759
thermoelectric 20 0.7714 0.8496 0.7987 0.9124
superconductor 56 0.9431 0.9180 0.8880 0.8228
Avg. - 0.7121 0.7804 0.7333 0.7743
@@ 4.4 Comparison with an existing approach
However, unlike CLaSP, which aligns structure–language pairs, CMML uses pairs of purely structural data (i.e., structure and XRD patterns). As a result, CMML produces abstract embeddings that limit human-intuitive understanding, and it also prevents a quantitative comparison in the text-based retrieval task (Sec. 4.1).
To compare how well these two approaches capture the high-level semantics of materials, we visually examined how they map semantically similar materials in their respective embedding spaces. Specifically, we generated embeddings of the crystal structures in the COD test set using both the CLaSP model and the CMML model (pre-trained on the Materials Project dataset) that is publicly provided by the authors. We then created t-SNE visualizations of these embeddings, highlighting entry points whose corresponding publication titles included specific keywords—specifically, ‘superconductor’ and ‘metal-organic framework.’
@@ 5.1 Use of titles and abstracts for supervision
A similar issue may explain the relatively low retrieval accuracy for the keyword ‘ferromagnetic’ in Fig. 3a and Table 1 (PT). Since ferromagnetism is a major characteristic of widely used iron-based materials, paper titles often omit the term ‘ferromagnetic.’ This could introduce noise into the title-based supervision and hinder the learning of this material concept. In contrast, Fig. 3b and and Table 1 (FT) show that fine-tuning using keywords derived from both titles and abstracts improves the overall retrieval performance. This suggests that abstracts convey more comprehensive information about the materials.
@@ 5.2 Dataset biases
Our analysis also revealed a potential bias in the COD dataset, which is predominantly composed of crystallography and chemistry publications. Over 80% of the entries come from a small subset of journals primarily focused on these fields (Fig. 7). This bias is understandable, given the COD’s historical development and its emphasis on crystallography. However, it highlights the need for more diverse data sources to ensure a comprehensive representation of materials and their properties. Given the limited availability of large materials databases with publication records beyond the COD, augmentation with external sources, such as citation contexts and Wikipedia entries, could be beneficial. This approach is similar to retrieval-augmented generation (RAG) [57] used in LLM applications. We leave such extensions for future work.
@@ 6 Conclusions and broader impacts
These promising results indicate the potential to transform how we explore the vast materials space. Potential applications include crystal structure screening and tagging via text prompts. Furthermore, inspired by text-to-image generation [58], CLaSP’s text embeddings could guide crystal structure generation models [59, 60], enabling innovative applications such as text-to-crystal generation. These advancements promise a more intuitive and efficient approach to exploring the materials space.
@@ 7 Method details
Below, we provide detailed procedures for model training, dataset creation, and embedding visualization.
@@ 7.3 Data preprocessing and spliting
We used the train set for title-based pre-training, the validation set for selecting a model checkpoint during pre-training or fine-tuning, and the test set for evaluating the retrieval performance (Figs. 3 and 4; Table 1) and for visualizing the embedding space (Figs. 5 and 6). We also used this dataset to generate the keyword-captioned dataset for fine-tuning, as explained next.
@@ 7.4 Keyword dataset generation for fine-tuning
The keyword generation process took 36 hours using a server equipped with eight NVIDIA A100 GPUs (80GB VRAM) and an efficient LLM inference framework, vLLM [65]. Finally, we removed generated keywords if they were unrelated to material properties, such as ‘crystal structure’, ‘X-ray diffraction’, ‘neutron diffraction’, ‘powder diffraction’, and ‘single-crystal X-ray diffraction’. Entries without any remaining keywords were also removed from the dataset.
@@ 7.5 Visualizations
Crystal structures in Figs. 1 and 2 were visualized using VESTA [66]. For the embedding space visualizations shown in Figs. 5 and 6, we employed the t-SNE algorithm [67] , as implemented in openTSNE [68]. In Fig.5, embedding clusters were generated using the k-means++ algorithm [69]. These clusters were further annotated by summarizing the associated paper titles using Google Gemini 1.5 Pro with a default temperature parameter of 1.0.

### v2

# Bridging Text and Crystal Structures: Literature-driven Contrastive Learning for Materials Science

@@ 1 Introduction
AI-driven materials science pursues this ambition through the use of machine learning (ML). One area of research has focused on predicting material properties using graph neural networks [3, 4, 5, 6] and transformers [7, 8, 9], leveraging large-scale crystal structure datasets annotated with properties simulated by first-principles calculations. Although this approach has shown success, these models are specialized for specific simulatable properties, such as bandgap, and are unable to provide a comprehensive view of materials with diverse properties and functionalities.
Other studies have explored developing embedding spaces for crystal structures to capture their similarities based on properties and functionalities [10, 11, 12, 13]. However, these efforts are bottlenecked by the lack of dedicated training datasets with diverse property and functionality annotations. Annotating crystal structures is costly, requiring expert knowledge, physical experimentation, or computationally intensive simulations. Consequently, these methods often produce abstract, unannotated embedding spaces that are not easily navigable for materials discovery. Because these spaces cannot be queried directly with natural-language descriptions, they remain of limited use to researchers who wish to search by desired properties rather than specific structural identifiers.
@@
To overcome this limitation, we introduce Contrastive Language–Structure Pre-training (CLaSP) (Fig. 1). CLaSP leverages a large-scale dataset of published crystal structures and corresponding article information retrieved from the Crystallography Open Database (COD) [23]. Specifically, we utilize publication titles for pre-training and keyword-based captions, generated from pairs of titles and abstracts using a large language model (LLM), for fine-tuning. We hypothesize that these textual sources provide a comprehensive representation of material characteristics. By leveraging bibliographic data as natural descriptions of structures, we bypass the need for labor-intensive, specialized annotations, enabling large-scale training of crossmodal models.
Our central objective is to learn a joint language-structure representation in which crystal structures are organized according to the semantic concepts expressed in human language. In this representation, a user can supply a free-form textual description (e.g., “narrow-bandgap material”) and retrieve candidate structures, while the model can simultaneously assign such semantic labels to previously unlabeled structures. This capability distinguishes our approach from conventional text-based search systems that rely solely on existing textual annotations. These conventional search systems falter when textual metadata are absent or incomplete for structures, which is frequently the case for newly simulated or experimentally determined structures.
We demonstrate the effectiveness of CLaSP through two key applications: intuitive text-based material retrieval and materials space mapping. Extensive analyses show that CLaSP effectively learns structure embeddings that capture abstract and complex material concepts, such as ‘superconductor’ and ‘metal-organic frameworks’.
@@ 2 Contrastive language–structure pre-training
To learn crystal embeddings $\{\bm{c}_{i}\}$, a CGCNN model [3] was trained from scratch with a redesigned output head. We replaced the original property regression layer with a linear projection that directly produces an embedding vector. Consequently, no property labels (e.g., formation energy) were used or predicted during training; the crystal encoder was supervised solely through the contrastive loss in Eq. (1) so that its embedding aligns with the paired text embedding. For text embeddings $\{\bm{t}_{i}\}$, a frozen pre-trained SciBERT model [25] was used as the text encoder, followed by a three-layer multilayer perceptron (MLP) fed with the CLS token embedding. We used the embedding dimensionality of 768 for both modalities. We provide the detailed training procedure in Sec. 7.1.
@@ 3.1 Crystal structure modeling
Crystal structure modeling using neural networks serves as a fundamental basis for ML models that process crystal structures. One of its most straightforward applications is crystal property prediction, for which various crystal encoder architectures have been proposed. These architectures typically employ geometric graph neural networks (GNNs) that model interatomic interactions via neural message passing. Early successful examples include CGCNN [3] and SchNet [31], which represent crystal structures as distance-based graphs. Subsequent work has aimed to capture richer interaction representations by incorporating bond and global state attributes [4], three-body interactions [32], attention mechanisms [33, 7, 8, 9], and infinite periodicity [6, 8].
@@ 3.2 Embedding learning for materials and crystal structures
To mitigate this limitation, more recent studies have explored extending material representations through multimodal learning with texts [21, 22], aiming to attribute semantic meaning to material embeddings. For example, Ozawa et al. [22] used crystal structures along with rule-based textual descriptions of their geometric features for contrastive learning. While the individual embedding dimensions of these structure embeddings are not meant to be interpreted directly, co-training with text in a shared latent space enables semantic interpretation via proximity to descriptive text embeddings—allowing for semantic inference even without manually labeled structures at deployment time. However, these methods rely on textual descriptions of structural features rather than material properties, thus limiting their ability to capture high-level information such as material functionalities.
@@ 4 Results
In this section, we demonstrate the effectiveness of the proposed approach both quantitatively and qualitatively through two applications: text-based material retrieval and embedding space visualization. For text-based material retrieval (Sec. 4.1), we evaluate how accurately the model links crystal structures with their textual descriptions, such as ‘superconductor’ and ‘narrow-bandgap material’. Additionally, we verify the design of our loss function through this retrieval application (Sec. 4.2). For embedding space visualization, we construct an intuitive map of the materials space (Sec. 4.3). We also utilize these two applications to compare the proposed method with an existing embedding approach (Sec. 4.4).
@@ 4.1 Zero-shot crystal structure screening by texts
To evaluate CLaSP’s ability to link crystal structures with textual property descriptions, we performed crystal structure retrieval using keywords representing material functionalities (e.g., ‘thermoelectric’ and ‘superconductor’). Given the embedding of a query keyword, we retrieved structure embeddings from the test set that showed high cosine similarities with the keyword embedding. Once trained, this application requires only a database of unannotated test structures, relying on publication-associated structures solely during training. By enabling semantic inference for previously unseen or unannotated structures, it represents a fundamental departure from conventional text-based retrieval systems.
For evaluation, we assessed the retrieval performance of the proposed approach by using publication titles associated with the test materials as hidden labels. We regarded a structure to possess a queried property if the associated paper title contained the keyword or its variations (e.g., given ‘superconductor’ as a query, the terms ‘superconductive’ and ‘superconductivity’ were also considered correct). This evaluation essentially demonstrates how competitively the proposed method performs compared to conventional text-based retrieval, despite the absence of textual annotations for the target structures.
Retrieval performance is reported with two complementary metrics. We examine the trade-off between true- and false-positive rates with the ROC (Receiver Operating Characteristic) curve. The area under this curve (ROC-AUC) equals the probability that a randomly chosen positive sample scores higher than a randomly chosen negative one, thus reflecting the overall ordering among all 40 604 candidates (test set). In contrast, Average Precision (AP) condenses the entire precision–recall curve into a single value by integrating precision over all recall levels. Thus, it emphasizes how densely true positives populate the upper part of the ranking—the first few dozen entries that most researchers actually inspect. Because our target materials account for less than 0.4% of the dataset (see Table 1), the overwhelming surplus of negatives would drive AP toward zero. To keep the AP scores informative, we evaluate AP on a subset of the test set where the negatives are randomly downsampled to match the number of positives. Overall, ROC-AUC measures the global discriminative power, whereas AP evaluates whether the positives are ranked near the top.
Table 1 summarizes the ROC-AUC and AP scores for six query keywords representing some functionality-level material concepts, evaluated using pre-trained (PT) and fine-tined (FT) models. Interestingly, even for material concepts with only a few hundred positive samples in the training set (corresponding to 0.05–0.38% of the training set), the model can still learn to identify various functionalities at a notably high performance level. Given the complexity of these material properties, this is quite remarkable and suggests that the learned representations can generalize effectively under limited supervision.
Figures 3a and 3b show detailed ROC curves for these six keywords before and after fine-tuning, with dashed diagonal lines representing random selection. While the zero-shot prediction using the pre-trained model (Fig. 3a) already demonstrates good performance, with a mean ROC-AUC of 0.7185, fine-tuning (Fig. 3b) further improved it to 0.7804. A similar trend is observed in the AP evaluation (Table 1).
@@
Positives (Train) ROC-AUC AP
Query # Positives (Test) # Count % Percent PT FT PT FT
ferromagnetic 165 1241 0.38% 0.4281 0.5674 0.5018 0.5716
ferroelectric 95 746 0.23% 0.5814 0.7097 0.5357 0.6380
semiconductor 90 719 0.22% 0.7382 0.8290 0.7705 0.8254
superconductor 56 465 0.14% 0.9431 0.9180 0.8880 0.8228
electroluminescence 24 217 0.07% 0.8106 0.8090 0.9050 0.8759
thermoelectric 20 161 0.05% 0.7714 0.8496 0.7987 0.9124
Avg. - - - 0.7121 0.7804 0.7333 0.7743
Figure 3: ROC curves of keyword-based crystal structure retrieval. (a) The zero-shot results with only pre-training show good performance, and (b) fine-tuning leads to further improvements. (c) The baseline method CMML [11], which is trained solely on crystal structure information and does not utilize any text descriptions, lags behind our proposed approach that leverages both structure and textual data. The test set consists of 40,604 materials, and the dashed diagonal lines represent the ROC curves for random selection.
@@ 4.3 Embedding space visualization
The matrix visualization in Fig. 6 further quantitatively analyzes the intra-cluster coherence and inter-cluster separation of the map shown in Fig. 5a. Each matrix element represents the ‘distance’ between two clusters, evaluated based on the Jensen–Shannon divergence (JSD) between histograms of words in paper titles (see Sec. 7.6 for for methodological details). In a well-formed clustering, the JSD values within clusters should be small, while those between different clusters should be large, resulting in a matrix heatmap with dark diagonal elements.
Despite the presence of noise in the JSD between title texts, Fig. 6 shows a prominent dark diagonal, indicating strong intra-cluster coherence and clear inter-cluster separation among the title groups.
We also observe two darker square blocks along the diagonal: one in the upper left and the other near the center. The first block corresponds to clusters dominated by minerals, intermetallic compounds, and oxides, while the second encompasses clusters rich in organometallic and coordination complexes. These block-wise proximities reflect the contiguous ‘lands’ seen in Fig. 5, showing that chemically related subfamilies are split into neighboring, finer-grained clusters that nonetheless remain close in the latent space.
@@
Figure 6: A Jensen–Shannon divergence matrix for cluster coherence and separation analysis. The rows and columns represent the 20 clusters from the map in Fig. 5a. Each matrix element represents the ‘distance’ between two clusters, evaluated based on the symmetric Jensen–Shannon divergence detailed in Sec. 7.6. Dark-colored diagonal elements indicate low intra-cluster divergence (i.e., high coherence), while bright-colored off-diagonal elements indicate high inter-cluster divergence (i.e., strong separation).
@@ 4.4 Comparison with an existing approach
However, CMML lacks a built-in mechanism for projecting text into its embedding space, thereby preventing direct comparison in text-based retrieval tasks. We therefore devised a proxy procedure that enables a fair, keyword-based retrieval comparison. For each keyword used in Sec. 4.1, we collected all COD training structures whose publication titles contained the keyword, averaged their CMML embeddings, and used the resulting vector as a representative concept embedding. Given this vector, retrieval proceeds exactly as with CLaSP: we compute the cosine similarities to the embeddings of the 40,604 COD test structures and rank them accordingly.
Figure 3c shows the ROC curves for CMML on the six keywords, in comparison to the CLaSP results shown in Fig. 3a. These results suggest that materials semantics cannot be captured without utilizing both structural and textual information.
To interpret these results, we next visually examined how each model maps semantically similar materials in its embedding space. We created t-SNE visualizations of these crystal embeddings, highlighting entry points whose corresponding publication titles included specific keywords—specifically, ‘superconductor’ and ‘metal-organic framework.’
@@ 5.1 Use of titles and abstracts for supervision
A similar issue may explain the relatively low retrieval accuracy for the keyword ‘ferromagnetic’ in Fig. 3a and Table 1 (PT). Since ferromagnetism is a major characteristic of widely used iron-based materials, paper titles often omit the term ‘ferromagnetic.’ This could introduce noise into the title-based supervision and hinder the learning of this material concept. In contrast, Fig. 3b and Table 1 (FT) show that fine-tuning using keywords derived from both titles and abstracts improves the overall retrieval performance. This suggests that abstracts convey more comprehensive information about the materials.
@@ 5.2 Dataset biases
Our analysis also revealed a potential bias in the COD dataset, which is predominantly composed of crystallography and chemistry publications. Over 80% of the entries come from a small subset of journals primarily focused on these fields (Fig. 8). This bias is understandable, given the COD’s historical development and its emphasis on crystallography. However, it highlights the need for more diverse data sources to ensure a comprehensive representation of materials and their properties. A promising alternative to the COD is the Inorganic Crystal Structure Database (ICSD), although it is distributed under a paid license and would prevent us from publicly releasing any resulting dataset if used. Given the limited availability of large materials databases with publication records beyond the COD and ICSD, augmentation with external sources, such as citation contexts and Wikipedia entries, could be beneficial. This approach is similar to retrieval-augmented generation (RAG) [58] used in LLM applications. We leave such extensions for future work.
@@
We also acknowledge a potential bias introduced by the LLM-generated keywords that we used for fine-tuning, as these keywords have not been formally validated by domain experts. Nevertheless, their use has demonstrably improved the performance of keyword-based materials retrieval (cf. Fig. 3a vs. 3b), which suggests that they possess a reasonable degree of semantic validity. A systematic expert evaluation of the generated keywords, along with iterative refinement of the prompting strategy used for keyword generation, remains an important direction for future work.
@@ 6 Conclusions and broader impacts
A fundamental finding of this study is that bibliographic metadata associated with crystal structures can effectively guide the learning of a crossmodal latent space that aligns structures with their intuitive textual descriptions. This was successfully demonstrated using simple publication metadata, specifically titles and abstracts. Given that the contrastive learning paradigm of CLaSP can flexibly incorporate text–structure pairs from any source, a natural and ambitious extension is to harness full-text materials science literature—a rich repository where decades of materials knowledge have been systematically articulated. In the era of data scaling laws [59], the field of materials science increasingly confronts the challenge posed by the scarcity of richly annotated materials data. The proposed literature-driven learning thus offers a realistic and scalable path forward for data-driven materials discovery.
The resulting crossmodal embedding not only underpins the retrieval and open-vocabulary classification shown here, but also paves the way for automatic captioning of crystal structures and for text-conditioned generation of novel candidates, echoing the evolution of vision–language models [60, 61]. Taken together, these capabilities promise a more intuitive and efficient exploration of the vast materials design space.
@@ 7 Method details
Below, we provide detailed procedures for model training, dataset creation, embedding visualization, and clustering quality analysis.
@@ 7.3 Data preprocessing and spliting
We used the train set for title-based pre-training, the validation set for selecting a model checkpoint during pre-training or fine-tuning, and the test set for evaluating the retrieval performance (Figs. 3 and 4; Table 1) and for visualizing the embedding space (Figs. 5 and 7). We also used this dataset to generate the keyword-captioned dataset for fine-tuning, as explained next.
@@ 7.4 Keyword dataset generation for fine-tuning
The keyword generation process took 36 hours using a server equipped with eight NVIDIA A100 GPUs (80GB VRAM) and an efficient LLM inference framework, vLLM [66]. Finally, we removed generated keywords if they were unrelated to material properties, such as ‘crystal structure’, ‘X-ray diffraction’, ‘neutron diffraction’, ‘powder diffraction’, and ‘single-crystal X-ray diffraction’. Entries without any remaining keywords were also removed from the dataset. Although this keyword generation step is time-consuming, it is performed only once in this study.
@@ 7.5 Visualizations
Crystal structures in Figs. 1 and 2 were visualized using VESTA [67]. For the embedding space visualizations shown in Figs. 5 and 7, we employed the t-SNE algorithm [68] , as implemented in openTSNE [69]. In Fig.5, embedding clusters were generated using the k-means++ algorithm [70]. These clusters were further annotated by summarizing the associated paper titles using Google Gemini 1.5 Pro with a default temperature parameter of 1.0.
7.6 Clustering quality analysis
Let each article title be represented by an L1-normalized TF-IDF (term frequency-inverse document frequency) [71] vector, denoted as $\bm{p}_{k}$. Each $\bm{p}_{k}$ can thus be interpreted as a histogram of the words in the title, normalized to form a probability distribution. The TF-IDF weighting scheme assigns lower weights to common, uninformative words such as “the” and “and” by incorporating their inverse frequency across the entire test set corpus.
Given a cluster $C_{i}$, we define its centroid as
$\bm{\mu}_{i}=\frac{1}{|C_{i}|}\sum_{k\in C_{i}}\bm{p}_{k}.$ (2)
We then compute the average Jensen–Shannon divergence between each cluster centroid and the entries in another cluster, resulting in the following matrix:
$M^{\mathrm{JS}}_{ij}=\frac{1}{|C_{j}|}\sum_{k\in C_{j}}JS\bigl{(}\bm{\mu}_{i},\,\bm{p}_{k}\bigr{)}.$ (3)
The visualization in Fig. 6 is based on the following symmetrized matrix:
$M^{\mathrm{SJS}}_{ij}=\frac{1}{2}\!\left(M^{\mathrm{JS}}_{ij}+M^{\mathrm{JS}}_{ji}\right).$ (4)
Intuitively, the diagonal elements of $M^{\mathrm{SJS}}_{jj}$ quantify intra-cluster coherence, with smaller values indicating tighter clusters. In contrast, the off-diagonal elements reflect inter-cluster separation, with larger values indicating better separation.